OpenStudy (anonymous):

Is there a difference between orthogonal basis and orthogonal complement of a subspace

OpenStudy (anonymous):

I need help finding the orthogonal subspace W of R5 spanned by vectors, V1,V2,V3,V4,V5

OpenStudy (anonymous):

I know how to find the orthogonal basis using the gram-schmidt process

OpenStudy (amistre64):

the orthogonal basis of a plane are 2 independant vectors in the plane whose dot product is 0; the orthogonal compliment is the setup such that every vector in the plane is orthogonal to a vector ... usually called the normal vector ... such that a(x-xo)+b(y-yo)+c(z-zo) = 0

OpenStudy (anonymous):

so one in r5 would be a(t-t0)+b(u-u0)+c(v-v0)+d(w-w0)+c(z-z0)=0

OpenStudy (amistre64):

that would be a suitable dot product yes :) but it might be simpler to say for higher dimensions, the we need to find the basis of the nullspace of A

OpenStudy (anonymous):

so AX=0 rref (A) * (x1,x2,x3,x4,x5)=0

OpenStudy (anonymous):

\[<-2,1,0,-1,-2>,<1,3,1,-2,-4>,<3,2,1,-1,-2>,<-1,4,1,-3,-6>,<-4,2,0,-2,-4>\]

OpenStudy (anonymous):

so after i find the null space of A would tht be my orthogonal component to the subspace W of r5

OpenStudy (amistre64):

the basis of the nullspace is the orthogonal compliment yes

OpenStudy (amistre64):

dunno if we would have to tweak the null basis tho to be orthogonal itself ... http://ltcconline.net/greenl/courses/203/Vectors/orthogonalComplements.htm

OpenStudy (anonymous):

i get (1,0,-1,1,2) (0,1,1,1,0)

OpenStudy (amistre64):

rref{[-2,1,3,-1,-4},{1,3,2,4,2},{0,1,1,1,0},{-1,-2,-1,-3,-2},{-2,-4,-2,-6,-4}} \[x_1=1x_3-1x_4-2x_5\\ x_2=-1x_3-1x_4+0x_5\\ x_3=1x_3+0x_4+0x_5\\ x_4=0x_3+1x_4+0x_5\\ x_5=0x_3+0x_4+1x_5\\\] \[\begin{pmatrix} 1&-1&-2\\ -1&-1&0\\ 1&0&0\\ 0&1&0\\ 0&0&1\\ \end{pmatrix}\]

OpenStudy (amistre64):

might have to dbl chk to see if i remembered that correctly

OpenStudy (anonymous):

well i got the same rref as the one you did, but did you then turn that into a system of equations and let x3,x4,x5 be variables and wrote everyting in terms of x1 and x2 and do parametric stuff

OpenStudy (amistre64):

yes i took and defined all the components of X in terms of the free variables

OpenStudy (amistre64):

there example at the bottom of the link used the vectors as row vectors

OpenStudy (amistre64):

nullspace of A^T rrefs to orthro comp of column vectors

OpenStudy (amistre64):

so maybe it should have gone this route rref{{-2,1,0,-1,-2},{1,3,1,-2,-4},{3,2,1,-1,-2},{-1,4,1,-3,-6},{-4,2,0,-2,-4}} http://www.wolframalpha.com/input/?i=rref%7B%7B-2%2C1%2C0%2C-1%2C-2%7D%2C%7B1%2C3%2C1%2C-2%2C-4%7D%2C%7B3%2C2%2C1%2C-1%2C-2%7D%2C%7B-1%2C4%2C1%2C-3%2C-6%7D%2C%7B-4%2C2%2C0%2C-2%2C-4%7D%7D <1,2,-7,0,0>, <1,-5,0,-7,0>,<2,-10,0,0,-7>

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how do u verify

OpenStudy (amistre64):

lol, thats a good question :) if we go back to the basics, how would you find the solution to Ax = 0?

OpenStudy (amistre64):

spose W = (a1,a2,a3,...,an); then W^T is such that a1 * x = 0, a2*x=0, a3*x=0, ..., an*x=0

OpenStudy (amistre64):

so it looks to me like if we take our vectors and represent them as the rows of the matrix A, then Ax = 0 represents the solution